Assume you are a soybean producer. On June 20, you decide to hedge the sale of a portion of your expected bean crop for delivery in the fall. Currently, November futures are trading at $7.20 per bushel. The next trading day, the November soybean contract settles at $6.95/bu. How much money is put into or taken out of your margin account by the clearing corporation? The initial and maintenance margin for soybean hedgers is $1000, and there are 5000 bushels per contract. Put your answer in terms of a positive (money put in) or a negative (money taken out) number.
Answer #1

November future are trading at Price $7.2 per bushel

We have 5000 bushels hedged in 1 future contract

Initial and maintenance margin are $1000

Therefore if price of contact falls and Total value of contract falls below margin then investor has to top up with some amount so that marginal requirements are met.

Earlier price of contract was $7.2/ bushel

Price of bushel other day is $6.95/ bushel

Hence loss in position is 7.2-6.95=0.25/bushel

Total loss in value=0.25*5000=1250

Therefore value of contract is fallen from 7.2*5000=36000 to 34750

But maintenance margin was 1000 therefore investor needed to maintain this margin he has to put jn 1250-1000=$250

Answer is $250